# On computational complexity and average-case hardness of shallow-depth Boson Sampling

### 06 Mar 2026 v4

@@ Appendix H On average-case hardness including collision outcomes
In Fig. 6, we present the average ratio of collision-free outcomes and standard deviation (error bar) for the photon number given by $N=4,8,12,16$, from the local random circuits over $(\mathcal{B}\mathcal{B}^{*})^{q}$ with the repetition number $q=1,2,3$ and global Haar random circuits. Remarkably, local random circuits for $(\mathcal{B}\mathcal{B}^{*})^{q}$ shows great resemblance to the global Haar random circuits: For all of the photon numbers $N=4,8,12,16$ and the repetition numbers $q=1,2,3$, local random circuits for $(\mathcal{B}\mathcal{B}^{*})^{q}$ exhibits almost identical average collision-free ratio (and also the standard deviation) to the global Haar random circuits. Given that the bosonic birthday paradox is guaranteed for the global Haar random circuit ensemble, also given that the mode number we used covers all recent large-sized Boson Sampling experiments [zhong2020quantum, zhong2021phase, madsen2022quantum, deng2023gaussian], our numerical result shows that the local random circuit ensemble for $(\mathcal{B}\mathcal{B}^{*})^{q}$ guarantees the bosonic birthday paradox at least for the near-term finite-sized experiments. Even though this does not guarantee the bosonic birthday paradox of the local random circuit ensemble for $(\mathcal{B}\mathcal{B}^{*})^{q}$ in the asymptotic limit as the system size scales, still, we believe that constant repetition number $q$ would be sufficient for the bosonic birthday paradox, due to the close resemblance to the global Haar random circuit ensemble as shown in the figure.
@@
Additionally, we similarly present a numerical simulation result in Fig. 6 for the mode number $M=32$ and photon number $N=4,8,12,16$, which indicates the saturated regime where the bosonic birthday paradox does not hold. Even in this regime, the collision patterns of local random circuits over $(\mathcal{B}\mathcal{B}^{*})^{q}$ also show a great resemblance to those of global Haar random circuits. This result gives hope that one can possibly extend our result to the saturated regime, by conducting a similar analysis for our circuit ensemble to the analysis for global Haar random circuits [bouland2023complexity].

### 11 Mar 2026 v5

@@ Appendix H On average-case hardness including collision outcomes
In Fig. 6, we present the average ratio of collision-free outcomes and standard deviation (error bar) for the photon number given by $N=4,8,12,16$, from the local random circuits over $(\mathcal{B}\mathcal{B}^{*})^{q}$ with the repetition number $q=1,2,3$ and global Haar random circuits. Remarkably, local random circuits for $(\mathcal{B}\mathcal{B}^{*})^{q}$ shows great resemblance to the global Haar random circuits: For all of the photon numbers $N=4,8,12,16$ and the repetition numbers $q=1,2,3$, local random circuits for $(\mathcal{B}\mathcal{B}^{*})^{q}$ exhibits almost identical average collision-free ratio (and also the standard deviation) to the global Haar random circuits. Given that the bosonic birthday paradox is guaranteed for the global Haar random circuit ensemble, also given that the mode number we used covers all recent large-sized Boson Sampling experiments [4, 5, 6, 7], our numerical result shows that the local random circuit ensemble for $(\mathcal{B}\mathcal{B}^{*})^{q}$ guarantees the bosonic birthday paradox at least for the near-term finite-sized experiments. Even though this does not guarantee the bosonic birthday paradox of the local random circuit ensemble for $(\mathcal{B}\mathcal{B}^{*})^{q}$ in the asymptotic limit as the system size scales, still, we believe that constant repetition number $q$ would be sufficient for the bosonic birthday paradox, due to the close resemblance to the global Haar random circuit ensemble as shown in the figure.
@@
Additionally, we similarly present a numerical simulation result in Fig. 6 for the mode number $M=32$ and photon number $N=4,8,12,16$, which indicates the saturated regime where the bosonic birthday paradox does not hold. Even in this regime, the collision patterns of local random circuits over $(\mathcal{B}\mathcal{B}^{*})^{q}$ also show a great resemblance to those of global Haar random circuits. This result gives hope that one can possibly extend our result to the saturated regime, by conducting a similar analysis for our circuit ensemble to the analysis for global Haar random circuits [29].